5.) A card game goes like this: You draw a card from a 52-card deck. If it is a face card (jack, queen or king), you win $5; otherwise, you lose $2. What is your expected value for this game? (Round your answer to one decimal place.)
Answer by stanbon(75887)

A card game goes like this: You draw a card from a 52-card deck. If it is a face card (jack, queen or king), you win $5; otherwise, you lose $2. What is your expected value for this game? (Round your answer to one decimal place.)
----------------
There are 12 face cards.
----
E(x) = (3/13)5 + (10/13)(-2) = (15-20)/13 = -7/13 = -53.8 cents
---------------
Cheers,
Stan H.
-------------